### v1

# AMBER - Advanced SegFormer for Multi-Band Image Segmentation: an application to Hyperspectral Imaging

Abstract
Deep learning has revolutionized the field of hyperspectral image (HSI) analysis, enabling the extraction of complex and hierarchical features. While convolutional neural networks (CNNs) have been the backbone of HSI classification, their limitations in capturing global contextual features have led to the exploration of Vision Transformers (ViTs). This paper introduces AMBER, an advanced SegFormer specifically designed for multi-band image segmentation. AMBER enhances the original SegFormer by incorporating three-dimensional convolutions to handle hyperspectral data. Our experiments, conducted on the Indian Pines, Pavia University, and PRISMA [^1] Project carried out using PRISMA Products, © of the Italian Space Agency (ASI), delivered under an ASI License to use. datasets, show that AMBER outperforms traditional CNN-based methods in terms of Overall Accuracy, Kappa coefficient, and Average Accuracy on the first two datasets, and achieves state-of-the-art performance on the PRISMA dataset.
keywords:
Hyperspectral Imaging, Vision Transformers, Remote Sensing
[…]
These authors contributed equally to this work.
[…]
These authors contributed equally to this work.
[…]
These authors contributed equally to this work.
[…]
These authors contributed equally to this work.
[…]
These authors contributed equally to this work.
[…]
These authors contributed equally to this work.
[…]
These authors contributed equally to this work.
[…]
These authors contributed equally to this work.
[…]
These authors contributed equally to this work.
[…]
These authors contributed equally to this work.
[…]
1 Introduction
Hyperspectral imaging (HSI) captures detailed spectral information across a wide range of wavelengths, offering a unique and rich data source for various remote sensing applications, including environmental monitoring, agriculture, and mineral exploration. The classification of hyperspectral images involves identifying the material composition of each pixel, a task that requires complex algorithms capable of handling the high dimensionality and complexity of the data.
In recent years, deep learning algorithms, particularly convolutional neural networks (CNNs), have become prominent in HSI classification due to their ability to learn hierarchical features from raw data. CNN-based models have shown considerable success in extracting spectral-spatial features. However, the inherent limitation of CNNs in capturing spectral-spatial long-range dependencies has spurred interest in transformer-based models.
Vision Transformers (ViTs)[1], with their self-attention mechanism, have demonstrated remarkable success in various image-processing tasks. They are adept at capturing global contextual information, which is crucial for HSI classification. Many deep learning models have adopted the transformer architecture for hyperspectral data, and have achieved good results.
This paper presents AMBER, an advanced version of SegFormer [2] tailored for hyperspectral image segmentation. By integrating three-dimensional convolutions, AMBER effectively processes multi-band images, preserving spatial dimensions and enhancing feature extraction. We evaluate AMBER on the Indian Pines, Pavia University, and PRISMA datasets, demonstrating its superior performance compared to traditional methods.
2 Related Works
Deep learning algorithms have transformed the landscape of hyperspectral image (HSI) pixel classification by learning complex and meaningful features hierarchically. Among these, convolutional neural networks (CNNs) have become a cornerstone for exploring spectral-spatial features within local windows, owing to their robustness and scalability ([3]).
In the last years, new CNN architectures have emerged with the goal of classifying/segmenting hyperspectral images. For example, Zhong, Li, Luo, and Chapman (2017) [4] introduced an end-to-end spectral-spatial residual network that employs residual blocks to continuously learn rich hyperspectral features. Paoletti, Haut, Fernandez-Beltran, et al. (2018) [5] proposed a pyramidal residual network that incrementally diversifies high-level spectral-spatial attributes across layers. Zhu, Jiao, Liu, Yang, and Wang (2020) [6] brought forth the residual spectral-spatial attention network (RSSAN), which cascades spectral and spatial attention modules for superior classification accuracy. Additionally, Chhapariya, Buddhiraju, and Kumar (2022) [7] designed a multi-level 3-D CNN with varied kernel sizes to extract multi-level features. Despite these advancements, CNNs still struggle to capture global contextual features and long-range dependencies for both hyperspectral and three-channel images due to their intrinsic locality.
The emergence of Vision Transformers (ViTs) is transforming image-processing tasks with their powerful modeling capabilities. Transformers have revolutionized natural language processing (NLP) and are now being increasingly used in various other fields, such as computer vision and hyperspectral imaging. The key innovation of the transformer model is the self-attention mechanism [8], which allows the model to weigh the importance of different input tokens (words or image patches) when generating an output. This enables the transformer to capture long-range dependencies and relationships within the data. Vision Transformers adapt the transformer architecture, originally designed for sequential data like text, to process image data. In ViTs, an image is divided into fixed-size patches, which are then linearly embedded and treated as input tokens for the transformer model. In this context, SegFormer [2] is an advanced vision transformer specifically designed for image segmentation tasks. It combines the strengths of transformers and CNNs, to effectively capture both global context and local details required for precise segmentation.
This approach was successfully adapted to hyperspectral imaging in many transformer-based models. In particular SpectralFormer (Hong et al., 2021) [9], achieving impressive classification performance. Building on this foundation, Zhang, Meng, Zhao, Liu, and Chang (2022) [6] emphasized the need for models that harness both global and local features. Sun, Zhao, Zheng, and Wu (2022) [10] developed the spectral-spatial feature tokenization transformer (SSFTT) to capture high-level semantic and spectral-spatial features, demonstrating excellent performance across multiple HSI datasets. Mei, Song, Ma, and Xu (2022) [11] proposed the group-aware hierarchical transformer (GAHT), which confines multi-head self-attention to local spatial-spectral contexts, thereby boosting classification accuracy. He et al. (2021) [12] suggested a spectral-spatial transformer combining a sophisticated CNN with a dense transformer, effectively capturing spatial features and spectral relationships, leading to robust classification results.
3 Methodology
This section introduces AMBER, a new and advanced SegFormer for Multi-Band Image Segmentation. As the original SegFormer [2], AMBER consists of two main modules: a hierarchical Transformer encoder which generates high-resolution coarse features and low-resolution fine features; and (2) a lightweight All-MLP decoder used to fuse multi-level features together. The final layer of the decoder performs dimensionality reduction to generate a bi-dimensional semantic segmentation mask from a three-dimensional input image. Unlike the original SegFormer, AMBER is suited to cope with multi-band images thanks to its three dimensional convolution. Moreover, AMBER preserves the input spatial dimensions $H$ and $W$: this in order to maximize accuracy at the price of alight increase in the number of trainable parameters. For an overview of the architecture, see Fig. 1.
Figure 1: The proposed AMBER framework consists of two main modules: A hierarchical Transformer encoder to extract spatial/spectral features; and a All-MLP decoder to fuse these multi-level features, reduce the spectral dimension, and predict the semantic segmentation mask.
Given an image of $D\times H\times W$ (where $D$ is the spectral dimension), we first divide it into patches of $3\times 3\times 3$ using a three-dimensional convolution layer (the patch size is the kernel dimension of the convolutional layer). We then use these patches as input to the hierarchical Transformer encoder to obtain multi-level features at 1/4, 1/8, 1/16 of the original image resolution. We then pass these multi-level features to the All-MLP decoder to predict the segmentation mask at a $\frac{D}{4}\times\frac{H}{4}\times\frac{W}{4}\times N_{cls}$ resolution, where $N_{cls}$ is the number of classes. To provide a detailed explanation of our model architecture, we will follow the same steps as in the original paper [2].
3.1 Hierarchical Transformer Encoder
We create a series of Mix Transformer encoders (MiT) specifically designed for semantic segmentation of multi-band images, such as hyperspectral images.
Hierarchical Feature Representation.
Given an input image with a resolution of $D\times H\times W$, we perform patch merging to obtain a hierarchical feature map $F_{i}$ with a resolution of $\frac{Dqw}{2^{i+1}}\times\frac{H}{2^{i+1}}\times\frac{W}{2^{i+1}}\times C_{i}$, where $i\in{1,2,3,4}$, and $C_{i+1}$ is larger than $C_{i}$.
Overlapped Patch Merging.
We utilize merging overlapping patches to avoid the need for positional encoding. To this end, we define $K$, $S$, and $P$, where $K$ is the three dimensional kernel size (or patch size), S is the stride between two adjacent patches, and $P$ is the padding size. Unlike the original SegFormer, in our experiments, we set $K=3$, $S=1$, $P=1$ ,and $K=3$, $S=2$, $P=2$ to perform overlapping patch merging. The patch size is intentionally kept small to preserve image details and avoid parameter explosion. $S=1$ preserves the original image spatial dimensions $H$ and $W$, avoiding the reduction of spatial dimensions by $1/4$.
Efficient Self-Attention.
In the multi-head self-attention process, each of the heads $Q,K,V$ have the same dimensions $N\times C$, where $N=D\times H\times W$ is the length of the sequence, the self-attention is estimated as:
$\text{Attention}(Q,K,V)=\text{Softmax}(\frac{QK^{T}}{\sqrt{d_{head}}})V.$ (1)
We utilize the same approach as in the original paper to simplify the self-attention mechanism from a complexity of $O(N^{2})$ to $O(\frac{N^{2}}{R})$. In our experiments, we set $R$ to $[64,16,4,1]$ from stage-1 to stage-4.
Mix-FFN.
We introduce Mix-FFN [2] which considers the effect of zero padding using a $3\times 3\times 3$ Conv in the feed-forward network (FFN). Mix-FFN can be formulated as:
$x_{out}=\text{MLP}(\text{GELU}(\text{Conv}_{3\times 3\times 3}(\text{MLP}(x_{in}))))+x_{in}$ (2)
where $x_{in}$ is the feature from the self-attention module. Mix-FFN mixes a $3\times 3\times 3$ convolution and an MLP into each FFN
3.2 Lightweight All-MLP Decoder
The All-MLP decoder consists of five main steps. First, multi-level features from the MiT encoder go through a MLP layer to unify the channel dimension. Then, in a second step, features are up-sampled to the original three-dimensional image and concatenated together. Third, a MLP layer is adopted to fuse the concatenated features. Fourth, a convolutional layer is used with a kernel $(D\times 1\times 1)$ to reduce the image dimension to a bi-dimensional prediction. Finally, a two-dimensional convolutional layer with a kernel size of 1 is used to predict the segmentation mask, which has a $H\times W\times N_{cls}$ resolution, where $N_{cls}$ is the number of classes.
4 Experimental and analysis
4.1 Data and experimental environment description
In order to test the proposed architecture, we utilize hyperspectral imaging, which records a detailed spectrum of light for each pixel and provides an invaluable source of information regarding the physical nature of the different materials, leading to the potential of a more accurate classification. Experiments were conducted firstly, on two of the most widely used public dataset - Indian Pines [13] and Pavia University [13] hyperspectral datasets - and, finally, on a hyperspectral image acquired from the PRISMA satellite. Here is an overview of the datasets.
Indian Pines and Pavia University datasets
The Indian Pines and Pavia University data sets consists of images with the absorption band caused by water mist removed. The Indian Pines image is obtained from an AVIRIS spectrometer in a forest area in Northwestern Indiana, USA. The image size is $145\times 145$ pixels, the wavelength is $0.40-2.50\,\mu$m, and the number of bands is $200$, the spatial resolution is $20$ m. The Pavia University image is obtained by ROSIS spectrometer at the University of Pavia, Italy. The image size is $610\times 340$ pixels, the wavelength is $0.43-0.86\,\mu$m, the number of bands is $103$, and the spatial resolution is $1.3$ m. The two datasets are divided into 16 categories and 9 categories respectively. Their false-color images and ground-truth images are shown in Figs. 2 and 3. The ground-truth classes and their respective samples number are shown in Table 1 and Table 2
Figure 2: Indian Pines dataset. (a) False-color map; (b) Ground-truth map.
Table 1: Groundtruth classes for the Indian Pines scene and their respective samples number.
# Class Samples
1 Alfalfa 46
2 Corn-notill 1428
3 Corn-mintill 830
4 Corn 237
5 Grass-pasture 483
6 Grass-trees 730
7 Grass-pasture-mowed 28
8 Hay-windrowed 478
9 Oats 20
10 Soybean-notill 972
11 Soybean-mintill 2455
12 Soybean-clean 593
13 Wheat 205
14 Woods 1265
15 Buildings-Grass-Trees-Drives 386
16 Stone-Steel-Towers 93
[…]
Figure 3: Pavia University dataset. (a) False-color map; (b) Ground-truth map.
Table 2: Groundtruth classes for the Pavia University scene and their respective samples number.
# Class Samples
1 Asphalt 6631
2 Meadows 18649
3 Gravel 2099
4 Trees 3064
5 Painted metal sheets 1345
6 Bare Soil 5029
7 Bitumen 1330
8 Self-Blocking Bricks 3682
9 Shadows 947
[…]
PRISMA dataset
PRISMA, a space-borne hyperspectral sensor developed by the Italian Space Agency (ASI), can capture images in a continuum of 231 spectral bands ranging between $400$ and $2500$ nm, at a spatial resolution of 30 m. The method was applied to a PRISMA Bottom-Of-Atmosphere (BOA) reflectance scene of the Quadrilátero Ferrífero area (Minas Gerias State, Brazil), located in the central portion of the Minas Gerais State (Northern Brazil).
The PRISMA hyperspectral image shows a reflectance scene at the Bottom-Of-Atmosphere (BOA) of the Quadrilátero Ferrífero mining district. It was acquired on May 21, 2022, at 13:05:52 UTC.
Level 2C PRISMA hyperspectral imagery was accessed from the mission website http://prisma.asi.it/. The VNIR and SWIR datacubes were pre-processed using tools available in ENVI 5.6.1 (L3Harris Technologies, USA). Errors in absolute geolocation were corrected using the Refine RPCs (HDF-EOS5) task, included within the PRISMA Toolkit in ENVI. Only 193 out of 231 bands were used for the final image due to atmospheric absorption affecting the excluded bands.
The ESA WorldCover-Map, available at https://esa-worldcover.org/en, was used in conjunction with the Global-scale mining polygons (Version 1) [14] to establish the ground truth for the Brazilian region. In order to avoid redundant and noisy information, only the following features were selected: vegetation (including dense, sparse, and very sparse vegetation as well as crops), water areas, mining areas, and build-up areas. As it can be seen from Table 3, the vegetation class in the image is over-represented. To address this issue, a random selection of $700,000$ vegetation pixels are set to 0 (undefined class). This adjustment ensures that the number of vegetation class pixels is comparable with the other classes. This is illustrated in Fig. 4.
Figure 4: PRISMA dataset. (a) False-color map; (b) Ground-truth map.
Table 3: Groundtruth classes for the PRISMA scene and their respective samples number.
# Class Samples
1 Vegetation 793572
2 Mining Area 36227
3 Permanent Water bodies 19638
4 Build-up 15463
[…]
4.2 Implementation details
We trained AMBER and made inferences using the IBiSCo Data Center (Infrastructure for Big Data and Scientific Computing). Information about the Data Center is available at https://ibiscohpc-wiki.scope.unina.it/. Specifically, AMBER was trained using four Tesla V100. Unlike the original SegFormer[2], training was performed from scratch without pre-training the encoder and randomly initializing the decoder. During the training phase, we applied random flipping to the patches and excluded the undefined class from the computation of the loss function. We set crop size to $D\times 32\times 32$ (where $D$ is the spectral dimension of the datasets) on all the three datasets. We trained the model using an SGD optimizer for all experiments. We used a batch size of 3 for experiments on the Indian Pines dataset, 10 for those on the Pavia University dataset, and 4 for those on the PRISMA dataset. In all of our experiments, we used the Categorical Cross Entropy loss function. The learning rate was set to 0.01 for all the experiments, and it remained constant throughout the training process. We used the following numbers of epochs for experiments on three datasets: 50 for Indian Pines, 30 for Pavia University, and 50 for PRISMA. We report semantic segmentation performance using Overall Accuracy (OA), Kappa Coefficient (Kappa), and Average Accuracy (AA). These metrics will be explained in the following section.
4.3 Evaluation metrics
In order to effectively assess the performance of this method in the experiment, three indicators, Overall Accuracy (OA), Kappa Coefficient (Kappa) and Avarage Accuracy (AA), are used to evaluate the classification performance of the model.
OA is calculated as the sum of correctly classified pixels divided by the total number of pixels. The number of correctly classified pixels is found along the diagonal of the confusion matrix, and the total number of pixels is equal to the total number of pixels of all real reference sources. The formula for OA is as follows:
$\text{OA}=\frac{\sum_{i=1}^{n}h_{ii}}{N}\times 100/\%$ (3)
Where:
  • $(h_{ii})$ is the number of correctly classified pixels distributed along the diagonal of the confusion matrix.
  • $N$ is the total number of samples.
  • $n$ is the number of categories.
The Kappa coefficient is a measure of consistency in testing. It is calculated by multiplying the total number $(N)$ of all real reference pixels by the sum of the diagonal $(h_{kk})$ of the confusion matrix. Then, the number of real reference pixels in each category and the classified pixels $(h_{ik},h_{jk})$ are subtracted from the product of the total number. This result is divided by the square of the total number of pixels minus the product of the total number of true reference pixels in each category and the total number of classified pixels in the category. The kappa coefficient comprehensively considers the various factors in the confusion matrix and provides a more comprehensive reflection of the accuracy of the overall classification. A larger value of the Kappa coefficient indicates higher accuracy of the corresponding classification algorithm. The general formula is as follows:
$\text{Kappa}=\frac{N\sum_{k}h_{kk}-\sum_{k}\left(\sum_{j}h_{kj}\sum_{i}h_{ik}\right)}{N^{2}-\sum_{k}\left(\sum_{j}h_{kj}\sum_{i}h_{ik}\right)}$ (4)
The Average Accuracy (AA) measures the average percentage of correctly classified samples for an individual class. It is calculated using the formula:
$\text{AA}=\frac{\sum_{i=1}^{C}\left({M_{ii}}/{\sum_{j=1}^{C}M_{ij}}\right)}{C}\times 100/\%$ (5)
Where $C$ is the total number of labels or classes, and $M_{ij}$ represents the samples that actually belonged to the ith class and were predicted to belong to the $j-th$ class.
4.4 Data Preprocessing
We constructed the datasets using only random patches, unlike other papers that employ various data balancing strategies (for example, see [15]). The purpose was both to assess AMBER’s ability to learn from an unbalanced dataset, and to create a dataset with a consistent number of patches. The center of each patch is a pixel randomly selected from pixels that are not located on the edges of the images. For the Pavia University dataset and the Indian Pines dataset, the center is a pixel that does not belong to the undefined class to avoid building patches with only undefined pixels. After creating the dataset, we addressed the well-known issue of pixel overlap between the training and test sets (effect of the random patches strategy), which can lead to information leakage [16, 17]. To mitigate this, we designated 20% of the patches for training and 80% for testing for the Indian Pines and Pavia University datasets, and 10% for training and 90% for testing for the PRISMA dataset.
Before the training phase, it is common to perform dimensionality reduction on the spectral dimension of hyperspectral images (e.g., PCA, SVD, or Gabor Filter [18, 19]) to address information redundancy. Conversely, our strategy was to keep the patches with the full spectral dimension, favoring a fine-grained classification, and leaving the task of extracting relevant features to AMBER’s encoder.
5 Results
5.1 Classification results
To evaluate the effectiveness of this method in classifying hyperspectral remote sensing images, we made a comparison with four mainstream methods: SVD/Unet [19], HSI-CNN [20], 3D-CNN [21] and contextual deep CNN [22]. It’s important to note that not all experiments involve pixel-by-pixel classification (or semantic segmentation). Only SVD/Unet, contextual deep CNN, and AMBER use this approach. On the other hand, HSI-CNN and 3D-CNN perform pure image classification, where the classes are determined by the centroid of the patch. This results in an increase in the metrics. Despite this, AMBER shows better results compared to the other methods on the Indian Pines dataset and Pavia University dataset and comparable results on the PRISMA dataset.
In the SVD/Unet method, the spectral dimension of the image is first reduced by a Singular Value Decomposition, considering only the first three singular values, and then a Unet architecture is used to classify pixels. This method has shown promising results in the semantic segmentation of PRISMA hyperspectral images.
In the HSI-CNN, the spectral-spatial features are first extracted from a target pixel and its neighbors. Next, several one-dimensional feature maps are obtained by performing a convolution operation on the spectral-spatial features, and these maps are then stacked into a two-dimensional matrix. Finally, this two-dimensional matrix, treated as an image, is input into a standard CNN.
The 3D-CNN architecture is a Multiscale 3D deep convolutional neural network (M3D-DCNN) which could jointly learn both 2D Multi-scale spatial feature and 1D spectral feature from HSI data in an end-to-end approach.
The contextual deep CNN first concurrently applies multiple 3-dimensional local convolutional filters with different sizes jointly exploiting spatial and spectral features of a hyperspectral image. The initial spatial and spectral feature maps obtained from applying the variable size convolutional filters are then combined together to form a joint spatio-spectral feature map. The joint feature map representing rich spectral and spatial properties of the hyperspectral image is then fed through fully convolutional layers that eventually predict the corresponding label of each pixel vector.
Every neural network mentioned above was trained on the same patches using the configuration described in the corresponding papers [19] [20] [21] [22]. “Same patches” in this context refers to patches that have identical centroids. AMBER was trained on patches of dimension $D\times 32\times 32$ where D is the spectral dimension, as previously mentioned. SVD/UNet was trained on patches of $3\times 32\times 32$ where 3 is the number of singular values. HSI-CNN was trained using patches of $D\times 3\times 3$, 3D-CNN using $D\times 7\times 7$ patches, and contextual deep CNN using $D\times 5\times 5$ patches (D is, again, the spectral dimension). For every method, we conducted 5 experiments for each dataset. The classification accuracy of different methods on the three data sets is presented in Tables 4, 5, and 6, with the highest values highlighted in bold font. A comprehensive comparative analysis is presented to draw conclusions.
Table 4: Classification accuracy of Indian Pines Dataset.
Class SVD/Unet [19] HSI-CNN [20] 3D-CNN [21] Cont Deep CNN [22] AMBER
1 $24.56$ $18.28$ $94.01$ $83.79$ $\boldsymbol{99.27}$
2 $\boldsymbol{99.70}$ $30.78$ $93.58$ $77.10$ $99.52$
3 $99.36$ $5.29$ $83.44$ $22.10$ $\boldsymbol{99.42}$
4 $57.24$ $7.66$ $77.85$ $48.69$ $\boldsymbol{99.49}$
5 $96.27$ $51.96$ $94.94$ $71.89$ $\boldsymbol{99.16}$
6 $99.96$ $96.09$ $99.69$ $99.25$ $\boldsymbol{99.98}$
7 $0.00$ $0.00$ $94.25$ $86.79$ $\boldsymbol{96.02}$
8 $91.98$ $95.30$ $99.63$ $93.49$ $\boldsymbol{99.97}$
9 $0.0029$ $1.56$ $93.65$ $34.85$ $\boldsymbol{96.95}$
10 $75.17$ $52.75$ $88.94$ $87.74$ $\boldsymbol{99.35}$
11 $99.93$ $89.80$ $95.83$ $89.77$ $\boldsymbol{99.86}$
12 $94.022$ $21.35$ $88.38$ $49.17$ $\boldsymbol{99.16}$
13 $73.19$ $97.36$ $\boldsymbol{100}$ $98.46$ $99.95$
14 $99.97$ $95.49$ $99.83$ $99.19$ $\boldsymbol{99.97}$
15 $72.0055$ $32.03$ $96.22$ $60.46$ $\boldsymbol{97.42}$
16 $73.58$ $79.07$ $98.94$ $99.90$ $\boldsymbol{99.54}$
OA/$94.18\pm 4.22$ $68.15\pm 7.36$ $95.01\pm 0.79$ $84.42\pm 2.02$ $\boldsymbol{99.74\pm 0.097}$
Kappa × 100 $92.77\pm 5.40$ $60.55\pm 10.029$ $94.03\pm 0.94$ $81.05\pm 2.56$ $\boldsymbol{99.68\pm 0.12}$
AA/$70.99\pm 4.22$ $47.059\pm 9.28$ $93.25\pm 1.10$ $71.40\pm 9.02$ $\boldsymbol{99.18\pm 0.34}$
[…]
Table 5: Classification accuracy of Pavia University Dataset.
Class SVD/Unet [19] HSI-CNN [20] 3D-CNN [21] Cont Deep CNN [22] AMBER
1 $89.19$ $96.79$ $98.94$ $97.48$ $\boldsymbol{99.79}$
2 $96.47$ $99.09$ $99.37$ $99.51$ $\boldsymbol{100}$
3 $0.17$ $74.44$ $90.70$ $82.06$ $\boldsymbol{99.27}$
4 $95.18$ $94.58$ $98.87$ $97.21$ $\boldsymbol{99.60}$
5 $94.15$ $99.89$ $\boldsymbol{100}$ $99.98$ $99.99$
6 $91.50$ $98.14$ $97.81$ $95.05$ $\boldsymbol{99.99}$
7 $44.09$ $96.90$ $97.06$ $97.12$ $\boldsymbol{99.88}$
8 $\boldsymbol{99.91}$ $96.34$ $98.29$ $96.05$ $99.85$
9 $99.75$ $99.70$ $99.86$ $99.96$ $\boldsymbol{99.98}$
OA/$91.12\pm 8.2$ $96.96\pm 0.77$ $98.51\pm 0.31$ $97.40\pm 0.84$ $\boldsymbol{99.94\pm 0.0098}$
Kappa × 100 $86.72\pm 11.82$ $96.15\pm 0.98$ $98.12\pm 0.39$ $96.49\pm 1.14$ $\boldsymbol{99.91\pm 0.015}$
AA/$70.99\pm 4.22$ $95.11\pm 1.66$ $97.87\pm 0.40$ $96.04\pm 0.64$ $\boldsymbol{99.82\pm 0.022}$
[…]
Table 6: Classification accuracy of PRISMA Dataset.
Class SVD/Unet [19] HSI-CNN [20] 3D-CNN [21] Cont Deep CNN [22] AMBER
1 $90.57$ $\boldsymbol{95.34}$ $93.59$ $77.97$ $81.70$
2 $94.99$ $68.97$ $81.75$ $89.52$ $\boldsymbol{96.60}$
3 $\boldsymbol{92.65}$ $74.32$ $78.69$ $87.80$ $92.21$
4 $\boldsymbol{96.71}$ $80.35$ $83.71$ $90.54$ $93.18$
OA/$\boldsymbol{93.52\pm 0.090}$ $89.88\pm 3.57$ $88.11\pm 0.21$ $88.52\pm 0.26$ $90.90\pm 0.84$
Kappa × 100 $\boldsymbol{91.06\pm 0.12}$ $74.62\pm 0.99$ $80.46\pm 0.51$ $84.06\pm 0.34$ $87.41\pm 1.48$
AA/$\boldsymbol{93.73\pm 0.10}$ $79.74\pm 0.69$ $84.44\pm 0.68$ $87.66\pm 0.54$ $90.92\pm 1.32$
[…]
Figure 5: Indian Pines prediction. (a) False-color map. (b) Ground-truth map. (c) AMBER prediction with the undefined mask. (d) AMBER prediction without the undefined mask.
Figure 6: Pavia University prediction. (a) False-color map. (b) Ground-truth map. (c) AMBER prediction with the undefined mask. (d) AMBER prediction without the undefined mask.
Figure 7: PRISMA prediction. (a) False-color map. (b) Ground-truth map. (c) AMBER prediction with the undefined mask. (d) AMBER prediction without the undefined mask.
The Overall Accuracy (OA), Kappa coefficients, and Average Accuracy (AA) of the AMBER method, calculated across 5 different experiments, are higher than those of the compared methods on both the Indian Pines dataset and the Pavia University dataset. On the Indian Pines dataset, the method achieved an OA of 99.74%, a Kappa coefficient of 0.9968, and an AA of 99.18%, representing a significant improvement compared to the other methods. In particular, AMBER demonstrates the ability to detect classes with few representations, such as class 7 (Grass-pasture-mowed), which has only 28 pixels. Other methods, such as SVD/Unet or HSI-CNN, struggle to detect this class, resulting in an accuracy per class of 0% . Another important feature is that, compared to other methods (especially SVD/Unet), AMBER has a very low standard deviation for OA, Kappa coefficient, and AA. This indicates a good ability to generalize: the results do not seem to depend on how the training data are chosen.
For the Pavia University dataset, the OA reached 99.94%, the Kappa coefficient reached 0.9991 and the AA reached 99.82%, mantaining an high accuracy on all the classes. As the Indian Pines dataset, AMBER has a very low standard deviation for OA, Kappa coefficient, and AA.
For the PRISMA dataset, the AMBER model achieved an Overall Accuracy of 90.90%, a Kappa coefficient of 0.8741, and an Average Accuracy of 90.92. While this result is slightly lower than the SVD/Unet method, it outperformed the other methods. Furthermore, the lower performances of AMBER on the PRISMA dataset are under investigation. AMBER is also able to improve the classification accuracy of a variety of features, such as Alfalfa, Grass-pasture-mowed and Soybean-mintill in the Indian Pines dataset; Asphalt, Gravel and Bitumen in the Pavia University dataset; mining area and build-up in the PRISMA dataset. This method is also able to perform segmentation on different types of hyperspectral images, maintaining high accuracy on both airborne (Pavia University and Indian Pines) and spaceborne (PRISMA) hyperspectral images.
The Fig. 5, 6, and 7 show the false color image (a), the ground truth sample plots (b), and the AMBER classification result plots (c) and (d). Specifically, plot (c) is the predicted result (d) with an additional mask of the ground truth undefined pixels applied a posteriori to visually justify the high accuracy for all types of features presented in Tables 4, 5 and 6. Specifically, on the Pavia University dataset and the Indian Pines dataset, the ground truth (b) and the prediction (c) are nearly identical, as expected.
Fig. 8 shows that AMBER accurately predicts the pixels classified as undefined in the ground-truth map (which are not used in the loss function computation). The figure shows an example of how AMBER can also segment complex structures. This particular structure belongs to the asphalt class in the Pavia University dataset.
Figure 8: AMBER prediction for the undefined pixels: a complex asphalt structure classification example
6 Conclusions
In this paper, we presented AMBER, an advanced SegFormer designed and fine tuned for hyperspectral image segmentation. By incorporating three-dimensional convolutions and preserving spatial dimensions, AMBER effectively captures both local and global features in multi-band images. Our extensive experiments on the Indian Pines, Pavia University, and PRISMA datasets show that AMBER in most cases outperforms traditional CNN-based methods and other state-of-the-art models in terms of Overall Accuracy, Kappa coefficient, and Average Accuracy.
The results indicate that AMBER not only improves classification accuracy but also exhibits robust generalization capabilities across different types of hyperspectral images. This demonstrates the potential of transformer-based architectures in advancing the field of hyperspectral image analysis. Future work could explore further optimizations and applications of AMBER in various remote sensing tasks, potentially leveraging additional data augmentation techniques and transfer learning strategies to enhance performance even further. Additionally, AMBER could be applied to astronomical data to extend its utility and effectiveness in analyzing complex and diverse datasets.
[…]
*Acknowledgements This work has been funded by project code PIR01_00011 “IBISCo”, PON 2014-2020, for all three entities (INFN, UNINA and CNR) and by the PNRR CN1 - ISP_S2_AGRINTESA for UNINA.